OpenStudy (loser66):

Calculate the linear transformation \(D^nS^n~~and~~S^nD^n,\) n= 1,2,3,.....on arbitrary element of P. D : differentiation; S: integration transformations

OpenStudy (loser66):

I know it is = P; just not know how to put in logic. D S =1 or DS =0? Please, help

OpenStudy (kinggeorge):

Is \(P\) the vector space of polynomials?

OpenStudy (loser66):

yes

OpenStudy (kinggeorge):

Well, my strategy would be to simply take an arbitrary polynomial of degree \(k\), and apply the two linear transformations on it. So for the transformation \(D^nS^n\), integrate the polynomial \(n\) times, and then take the derivative \(n\) times. Since everything is a polynomial, this shouldn't be too bad.

OpenStudy (loser66):

linear transformation D^n S^n = (DS)^n , is it right?

OpenStudy (kinggeorge):

For that particular product of transformation I think that's true, but it's not true in general. For example, \(S^nD^n\neq(SD)^n\).

OpenStudy (loser66):

I don't think so. Let say D*D*D*D*..........DS*S*S*S.........S= DS at the end

OpenStudy (loser66):

And you are right if the transformations do not cancel out like this D and S

OpenStudy (loser66):

So?? in this particular case, what should I conclude?

OpenStudy (kinggeorge):

I still think what you should do is just take an arbitrary polynomial, and apply the linear transformations. So let \[p(x)=a_kx^k+a_{k-1}x^{k-1}+...+a_1x+a_0\]and apply \(D^nS^n\) and \(S^nD^n\). One note for \(S^nD^n\) though, you need to consider the case when \(k\ge n\) and when \(k<n\) separately.

OpenStudy (loser66):

Ok, let do S^nD^n first, I have to take differentiate of p(x) n times. if n>k, at the time n = k_i I get a constant, then n = k_i +1 , I get 0. the case is done.

OpenStudy (kinggeorge):

Hold on. If \(n>k\), then \(D^n(p(x))=0\). But \(S^n(0)=g(x)\) where \(g(x)\) is a polynomial of degree \(n-1\).

OpenStudy (loser66):

Let make the simplest case S^3D^3(x+1) = S^3 D^2(2)=S^3D(0) = S^3 (0) no more polynomial to take integral, right? this 0 is a number 0, not polynomial degree 0

OpenStudy (loser66):

*S^3D^2(1)

OpenStudy (kinggeorge):

In this particular vector space, 0 is the zero vector, which is the zero polynomial.

OpenStudy (loser66):

I 'm not with you then. p(x) = \(\sum_k a_kx^k\) and \(a_k \) is a scalar. so, if the n>k, the term n=k is a number on D function

OpenStudy (kinggeorge):

Even if you do just consider 0 as a number, then \(\int0\,dx=a\) where \(a\in\mathbb{R}\) is a constant

OpenStudy (loser66):

and then , at the end up, how can we compute the degree of polynomial? If n is >>k, like p(x) = x+1, after twice, we get 0 for D, and start S. after twice we get x+ c ( another scalar, not 1 anymore) 3rd round = x^2 /2 + cx + b 4th round = x^3/6 +cx^2/2 + bx + d...... can it go on that way?

OpenStudy (kinggeorge):

What you should eventually see, is that if \(n>k\), is that \(S^nD^n(p(x))=g(x)\) where \(g(x)\) is an arbitrary polynomial of degree \(n-1\).

OpenStudy (loser66):

I got you :) I will try that way. thank you